Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufacturing departments--Molding and Fabrication. It started, completed, and sold only two jobs during March—Job P and Job Q. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March):

Molding Fabrication Total
Estimated total machine-hours used 2,500 1,500 4,000
Estimated total fixed manufacturing overhead $ 11,750 $ 16,050 $ 27,800
Estimated variable manufacturing overhead per machine-hour $ 2.10 $ 2.90
Job P Job Q
Direct materials $ 20,000 $ 11,500
Direct labor cost $ 26,600 $ 10,300
Actual machine-hours used:
Molding 2,400 1,500
Fabrication 1,300 1,600
Total 3,700 3,100

Sweeten Company had no underapplied or overapplied manufacturing overhead costs during the month.

Required:

For questions 1-8, assume that Sweeten Company uses a plantwide predetermined overhead rate with machine-hours as the allocation base. For questions 9-15, assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments.

1. What was the company’s plantwide predetermined overhead rate? (Round your answer to 2 decimal places.)

2. How much manufacturing overhead was applied to Job P and how much was applied to Job Q? (Do not round intermediate calculations.)

3. What was the total manufacturing cost assigned to Job P? (Do not round intermediate calculations.)

4. If Job P included 20 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)

5.  What was the total manufacturing cost assigned to Job Q? (Do not round intermediate calculations.)

6. If Job Q included 30 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)

7. Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establish selling prices for all of its jobs. What selling price would the company have established for Jobs P and Q? What are the selling prices for both jobs when stated on a per unit basis assuming 20 units were produced for Job P and 30 units were produced for Job Q? (Do not round intermediate calculations. Round your final answers to nearest whole dollar.)

8. What was Sweeten Company’s cost of goods sold for March? (Do not round intermediate calculations.)

For questions 9-15, assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments.

9. What were the company’s predetermined overhead rates in the Molding Department and the Fabrication Department? (Round your answers to 2 decimal places.)

10. How much manufacturing overhead was applied from the Molding Department to Job P and how much was applied to Job Q? (Do not round intermediate calculations.)

11. How much manufacturing overhead was applied from the Fabrication Department to Job P and how much was applied to Job Q? (Do not round intermediate calculations.)

12. If Job P included 20 units, what was its unit product cost? (Do not round intermediate calculations.)

13. If Job Q included 30 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)

14. Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establish selling prices for all of its jobs. What selling price would the company have established for Jobs P and Q? What are the selling prices for both jobs when stated on a per unit basis assuming 20 units were produced for Job P and 30 units were produced for Job Q? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)

15. What was Sweeten Company’s cost of goods sold for March? (Do not round intermediate calculations.)
Answer #1

(1) Company’s plantwide predetermined overhead rate :-

Estimated Total Manufacturing Overhead / Total machine hours

Estimated total overhead = Estimated total fixed manufacturing overhead (27800) + variable overhead molding(machine hours:2500*var overhead per machine hour molding:2.10) + variable overhead fabrication (machinehours:1500*var overhead per machine hour fabrication:2.90) = 26200 + 5250 + 4350 = 35800

$35800/4000 hours = $8.95 per machine hour

(2) Manufacturing overhead = (predetermine overheadrate*machine hours)

Job P = overhead rate * total machine hours = 8.95 (previous solution) * 3700 = $33115

Job Q= overhead rate * total machine hours = 8.55 (previous solution) * 3100 = $26505

(3) Total manufacturing cost = direct materials + direct labor + manufacturing overhead(previous solution) = 20000 + 26600 + 33115 = $79715